# On the Symbol Length of Fields with finite Square Class Number

### v3

@@ Abstract.
Let $F$ be a field of characteristic not $2$ with finitely many square classes. Using combinatorial arguments applied to objects related to vector spaces over finite fields, we deduce an upper bound for the number of Pfister forms over $F$. Moreover, we compute upper bounds for the $n$-symbol length $F$ ($n\in\operatorname{\mathbb{N}}$), i.e., the smallest integer $\operatorname{\mathrm{sl}}_{n}(F)\geq 0$ such that to each quadratic form $\phi\in\operatorname{I}^{n}(F)$ there exists some $0\leq k\leq\operatorname{\mathrm{sl}}_{n}(F)$ and Pfister forms $\pi_{1},\ldots,\pi_{k}$ such that $\varphi\equiv\pi_{1}+\ldots+\pi_{k}\mod\operatorname{I}^{n+1}(F)$. In particular, we rediscover a bound that can also be deduced from a result by Bruno Kahn that he stated without proof.
Key words and phrases:
quadratic form, Pfister forms, symbol length, Milnor $K$-theory, Galois cohomology, level, Pythagoras number.
1991 Mathematics Subject Classification:
11E04, 11E81, 12D15, 12G05, 19D45
@@ 1. Introduction
The determination of the symbol length is an active field of research, see, e.g., [Sal15], [Mat16] or [Cha20] for some recent progress in the field. The $3$-symbol lengths (and the related concept of $3$-Pfister number. i.e., the minimal length of a representation of forms in $I^{3}(F)$ as sums of general $3$-fold Pfistter forms) has, for example, been studied in [BRV10], [PST09], [Rac13], [Lor23].
The aim of this article is to find upper bounds for $\operatorname{\mathrm{sl}}_{n}(\varphi)$ and $\operatorname{\mathrm{sl}}_{n}(F)$ in the case where the square class group $F^{*}/F^{*2}$ is finite, i.e., $q(F)=|F^{*}/F^{*2}|=2^{d}<\infty$. As this problem is rather trivial for $n=1$ and in order to avoid subtleties in our calculations, we assume throughout that $n$ be an integer $\geq 2$. Note that K. Becher and the first author obtained best possible general upper bounds in the case $n=2$ in terms of the cardinality $q(F)=|F^{*}/F^{*2}|=2^{d}$, see [BH04].
@@
In Section 2, we provide a counting argument for the number of Pfister forms of a certain type over a field with finitely many square classes, and provide upper bounds for the size of certain quotient groups in a filtration of the square class group of a field with finitely many square classes. Using these results, we use more refined counting methods to obtain bounds for the $n$-symbol length (Theorem 3.8). The bound is of a rather technical nature and can be weakend to yield a polynomial bound in $d$ of degree $n$ (Corollary 3.10). Such a polynomial can be computed explicitly (see Example 3.12 where such polynomials are explicitly computed in the case $n=3$). A polynomial bound of that type can also be deduced from a result that was stated by B. Kahn but without proof, [Kah05, Proposition 2.3(h)].
@@
Isometry of two forms $\varphi_{1},\varphi_{2}$ will be denoted by $\varphi_{1}\cong\varphi_{2}$, their orthogonal sum by $\phi_{1}\perp\phi_{2}$, and their tensor product by $\phi_{1}\otimes\phi_{2}$. The orthogonal sum of $m$ copies of $\phi$ will be written $m\times\phi$. By abuse of notation, we will denote the Witt class of a quadratic form $\varphi$ in the Witt ring $WF$ again by $\varphi$.
@@
The level of $F$, denoted by $s(F)$, is the least integer $k$ such that $-1$ is a sum of $k$ squares in $F$, or infinity if no such $k$ exists. A famous result by A. Pfister [Pfi65] states that $s(F)$ is a power of $2$ if finite, and that each power of $2$ occurs as a level of a suitable field. A field is called formally real (or real for short), if $-1$ is not a sum of squares, and nonreal otherwise. The Pythagoras number of $F$, denoted by $p(F)$ is the least integer $k$ such that any sum of squares in $F$ is a sum of $k$ squares, or $\infty$ if no such integer exists. For nonreal fields, we have $s(F)\leq p(F)\leq s(F)+1$, see [Lam05, Chapter XI. Theorem 5.6]. For real fields, it was shown by the first author in [Hof99, Theorem 1] that every positive integer can be realized as the Pythagoras number of a suitable field.
@@
We will use several combinatorial arguments in the sequel. We need the binomial coefficient $\binom{n}{k}$ for integers $n,k\in\operatorname{\mathbb{Z}}$, i.e., the number of subsets witgh $k$ elements of a set with $n$ elements. For $0\leq k\leq n$, we have $\binom{n}{k}=\frac{n!}{k!(n-k)!}$ and in all other cases, we have $\binom{n}{k}=0$. Finally, for a set $S$, we denote the power set of $S$, i.e., the set consisting of all subsets of $S$ by $\mathcal{P}(S)$. More generally, for $m\in\operatorname{\mathbb{N}}$, we denote by $\mathcal{P}_{m}(S)$ the set of all subsets $A\subseteq S$ with $|A|=m$. In particular, if $S$ is finite, then $|\mathcal{P}_{m}(S)|=\binom{|S|}{m}$.
@@ 2. The Number of Pfister Forms
[…]
@@
Furthermore, any basis from a subspace $U$ can be transformed into any other basis of $U$ by successivly replacing a subset $\{x,y\}$ by $\{y,x\}$ or by $\{x,xy\}$. Also, we have isometries $\langle\!\langle-x,-xy\rangle\!\rangle\cong\langle\!\langle-x,-y\rangle\!\rangle\cong\langle\!\langle-y,-x\rangle\!\rangle$ for all $x,y\in F^{\ast}$. The map is thus well-defined and we have proved (a). Finally, for (b), every form in $\mathsf{P}_{n,m}(F)$ has a preimage due to Lemma 2.4. ∎
@@
In [Cor73], Cordes introduced so-called $\overline{C}$-fields. These are fields with finite square class group, i.e., $q(F)<\infty$, and such that $|\mathsf{W}(F)|=2^{q(F)}$. Such fields will always have level $s(F)=1$ or $s(F)=2$. Over such fields, every subset of $F^{\ast}/F^{\ast 2}$ is the value set of a unique anisotropic quadratic form over $F$ by [Lam05, Chapter XI. Theorem 7.19 (4)]. In particular, we have $\mathsf{D}_{F}(\pi_{U,0})=U$ and $\mathsf{D}_{F}(\pi_{U,1})=U\cup-U$. One now readily sees that each form in $\mathsf{P}_{n,m}(F)$ has exactly one preimage under $\varphi_{n,m}$. Note that $\overline{C}$-fields are examples of so-called rigid fields whose Pfister number, an invariant related to the symbol length, was studied by the second author in [Lor23].
@@
It is easy to calculate $s_{m}$ in terms of the level $s(F)$.
@@
Let $0\leq m\leq n$ be integers and let $F$ be a field with $d_{m}<\infty$. Suppose furthermore that $p(F)$ is finite, i.e., $2^{s}\leq p(F)<2^{s+1}$ for some integer $s\geq 0$. Then
$|\mathsf{P}_{n,m}(F)|\leq 2^{(n-m)(d_{m}+s_{m}-n+m)}\cdot\alpha_{m},$
where $\alpha_{m}$ is defined as in Corollary 2.2.
@@
The rest of the section is now devoted to find upper bounds for the $d_{m}$ in order to optimize the above bound for the number of Pfister forms. We will see in the next result that we can express $d_{m}$ in terms of $s_{m}$ and the $q_{k}$ for $k\leq m$.
@@
We now want to collect estimates for the invariants $q_{m}$.
To do so we will use the following extension of a classical result from I. Kaplansky.
@@
Let $F$ be a field with $q(F)=2^{d}<\infty$ and with finite $p(F)$, i.e., $2^{s}\leq p(F)<2^{s+1}$ for some integer $s\geq 0$.
@@ 3. Bounding the Symbol Length
$\mathcal{B}\supseteq\mathcal{B}_{0}\supseteq\mathcal{B}_{1}\supseteq\mathcal{B}_{2}\supseteq\ldots$
@@
If we have $\pi_{j}\in\mathsf{P}_{n,m}(F)$ for some $m\in\operatorname{\mathbb{N}}$, the resulting Pfister forms are hyperbolic or will lie in $\mathsf{P}_{n,m^{\prime}}(F)$ for some $m^{\prime}\geq m$. We can clearly omit the hyperbolic Pfister forms in our representation and thus now consider any of the resulting Pfister forms lying in $\mathsf{P}_{n,m^{\prime}}(F)$ for some $m^{\prime}\geq m$ If we have $m^{\prime}=m$, we can repeat the above until the considered slot in any newly introduced Pfister form lies in $\mathcal{B}_{m}$. If we have $m^{\prime}>m$, we first substitute the representation of this form with an isometric Pfister form with $m^{\prime}$ slots equal to $-1$ and all other slots being products of $-1$ and of elements lying in $\mathcal{B}_{m^{\prime}}$. We can then apply the above procedure to this new form. By repeating this process for all of the $\pi_{j},j\in\{1,\ldots,k\}$ and all slots that neither lie in the appropriate $\mathcal{B}_{m}$ (up to a factor $-1$) nor are equal to $-1$, and possibly rearranging the terms, we will eventually obtain a representation as desired. ∎
@@
If $F$ is nonreal of level $s(F)=2^{s}$, it is obviously enough to consider the sum just going up to $\min\{s,n\}$.
@@
as in Proposition 3.4. Let $X\in\mathcal{P}_{m,r}$ and $Y\in\mathcal{P}_{m,r}^{\prime}$. Then all forms $\pi\in\mathsf{P}_{n,m}(F)\cup\mathsf{P}_{n,m+1}(F)$ that occur in this representation and that have the elements of $X\cup Y$ as slots (up to a factor $-1$) can be replaced by a single Pfister form using Lemma 3.3 just as in the proof of Corollary 3.7. Conversely, let some $\pi=\pi_{U,m}$ from this representation be given and let $B=\{-b_{1},\ldots,-b_{n-m}\}$ be the set of slots $\neq-1$ of $\pi$. Let $r$ be maximal such that there is an $X\in\mathcal{P}_{m,r}$ with $X\subseteq-B$. We clearly have
@@
There is thus a subset $Y\subseteq\mathcal{A}_{m}^{\prime}$ with cardinality $|Y|=n-m-2r-1$ such that $X\cup Y\subseteq-B$. Hence, for all Pfister forms of type $\pi_{U,m}$ for a suitable $U$, there is an $r\in\operatorname{\mathbb{N}}$, $X\in\mathcal{P}_{m,r}$ and $Y\in\mathcal{P}_{m,r}^{\prime}$ such that all slots except one are either $-1$ or lie (up to a factor $-1$) in $X\cup Y$. Therefore, our above strategy to replace certain Pfister forms by a single one can be applied to all forms on the right hand side in (1). The symbol length $\operatorname{\mathrm{sl}}_{n}(\varphi)$ can thus be bounded by the cardinality of
@@
We now close the article with a comparison of these bounds for small level and Pythagoras number $p(F)=s(F)=2^{s}$ with $s\in\{0,1,2\}$. Before giving some explicit numbers, we want to recall the upper bounds that we found above.
First of all, we want to collect upper bounds for $d_{m}$ for some cases. In the following table, we have summarized the upper bounds we know for $d_{m}$ obtained in Corollary 2.13: $m$ $s$ 0 1 2 0 $d$ $d-1$ $d-1$ 1 00 $d-1$ $d-3$ 2 00 00 $d-3$ For $s_{m}$ we have the following values according to Remark 2.8 (a): $m$ $s$ 0 1 2 0 00 $1$ $1$ 1 00 00 $1$ 2 00 00 00

### v4

# On the Number of Pfister Forms and the Symbol Length of Fields With Finite Square Class Number

@@ Abstract.
Let $F$ be a field of characteristic not $2$ with finitely many square classes. Using combinatorial arguments applied to objects related to vector spaces over finite fields, we deduce an upper bound for the number of Pfister forms over $F$. Moreover, we compute upper bounds for the $n$-symbol length of $F$ ($n\in\operatorname{\mathbb{N}}$), i.e., the smallest integer $\operatorname{\mathrm{sl}}_{n}(F)\geq 0$ such that to each quadratic form $\phi\in\operatorname{I}^{n}(F)$ there exists some $0\leq k\leq\operatorname{\mathrm{sl}}_{n}(F)$ and Pfister forms $\pi_{1},\ldots,\pi_{k}$ such that $\varphi\equiv\pi_{1}+\ldots+\pi_{k}\mod\operatorname{I}^{n+1}(F)$. In particular, we rediscover a bound that can also be deduced from a result by Bruno Kahn that he stated without proof.
Classification (MSC 2020): 11E04, 11E81, 12D15, 12G05, 19D45
Keywords: quadratic form, Pfister forms, symbol length, Milnor $K$-theory, Galois cohomology, level, Pythagoras number
@@ 1. Introduction
The determination of the symbol length is an active field of research, see, e.g., [Sal15], [Mat16] or [Cha20] for some recent progress in the field. The $3$-symbol length (and the related concept of $3$-Pfister number, i.e., the minimal length of a representation of forms in $I^{3}(F)$ as sums of general $3$-fold Pfistter forms) has, for example, been studied in [BRV10], [PST09], [Rac13], [Lor23].
The aim of this article is to find upper bounds for $\operatorname{\mathrm{sl}}_{n}(\varphi)$ and $\operatorname{\mathrm{sl}}_{n}(F)$ in the case where the square class group $F^{*}/F^{*2}$ is finite, i.e., $q(F)=|F^{*}/F^{*2}|=2^{d}<\infty$. As this problem is rather trivial for $n=1$ and in order to avoid subtleties in our calculations, we assume throughout that $n$ be an integer $\geq 2$. Note that K. Becher and the first author obtained best possible general upper bounds in the case $n=2$ in terms of the cardinality $q(F)$, see [BH04].
@@
In Section 2, we provide a counting argument for the number of Pfister forms of a certain type over a field with finitely many square classes, and provide upper bounds for the size of certain quotient groups in a filtration of the square class group of a field with finitely many square classes. Using these results, we use more refined counting methods to obtain bounds for the $n$-symbol length (Theorem 3.8). The bound is of a rather technical nature and can be weakened to yield a polynomial bound in $d$ of degree $n-1$ (Corollary 3.10). Such a polynomial can be computed explicitly (see Example 3.12 where such polynomials are explicitly computed in the case $n=3$). A polynomial bound of that type can also be deduced from a result that was stated by B. Kahn but without proof, [Kah05, Proposition 2.3(h)].
@@
Isometry of two forms $\varphi_{1},\varphi_{2}$ will be denoted by $\varphi_{1}\cong\varphi_{2}$, their orthogonal sum by $\phi_{1}\perp\phi_{2}$, and their tensor product by $\phi_{1}\otimes\phi_{2}$. The orthogonal sum of $m$ copies of $\phi$ will be written $m\times\phi$. By abuse of notation, we will denote the Witt class of a quadratic form $\varphi$ in the Witt ring $\mathsf{W}(F)$ again by $\varphi$.
@@
The level of $F$, denoted by $s(F)$, is the least integer $k$ such that $-1$ is a sum of $k$ squares in $F$, or $\infty$ if no such $k$ exists. A famous result by A. Pfister [Pfi65] states that $s(F)$ is a power of $2$ if finite, and that each power of $2$ occurs as a level of a suitable field. A field is called formally real (or real for short), if $-1$ is not a sum of squares, and nonreal otherwise. The Pythagoras number of $F$, denoted by $p(F)$ is the least integer $k$ such that any sum of squares in $F$ is a sum of $k$ squares, or $\infty$ if no such integer exists. For nonreal fields, we have $s(F)\leq p(F)\leq s(F)+1$, see [Lam05, Chapter XI. Theorem 5.6]. For real fields, it was shown by the first author in [Hof99, Theorem 1] that every positive integer can be realized as the Pythagoras number of a suitable field.
@@
We will use several combinatorial arguments in the sequel. We need the binomial coefficient $\binom{n}{k}$ for integers $n,k\in\operatorname{\mathbb{Z}}$, i.e., the number of subsets with $k$ elements of a set with $n$ elements. For $0\leq k\leq n$, we have $\binom{n}{k}=\frac{n!}{k!(n-k)!}$ and in all other cases, we have $\binom{n}{k}=0$. Finally, for a set $S$, we denote the power set of $S$, i.e., the set consisting of all subsets of $S$ by $\mathcal{P}(S)$. More generally, for $m\in\operatorname{\mathbb{N}}$, we denote by $\mathcal{P}_{m}(S)$ the set of all subsets $A\subseteq S$ with $|A|=m$. In particular, if $S$ is finite, then $|\mathcal{P}_{m}(S)|=\binom{|S|}{m}$.
@@ 2. The Number of Pfister Forms
$V(2,d,m)\leq 2^{m(d-m)}\alpha_{m}\quad\mbox{where}\quad\alpha_{m}=\prod_{\ell=1}^{m}\left(1+\frac{1}{2^{\ell}-1}\right).$
@@
Furthermore, any basis from a subspace $U$ can be transformed into any other basis of $U$ by successively replacing a subset $\{x,y\}$ by $\{y,x\}$ or by $\{x,xy\}$. Also, we have isometries $\langle\!\langle-x,-xy\rangle\!\rangle\cong\langle\!\langle-x,-y\rangle\!\rangle\cong\langle\!\langle-y,-x\rangle\!\rangle$ for all $x,y\in F^{\ast}$.
The map is thus well-defined and we have proved (a). Finally, for (b), every form in $\mathsf{P}_{n,m}(F)$ has a preimage due to Lemma 2.4. ∎
@@
In [Cor73], Cordes introduced so-called $\overline{C}$-fields. These are fields with finite square class group, i.e., $q(F)<\infty$, and such that $|\mathsf{W}(F)|=2^{q(F)}$. Such fields will always have level $s(F)=1$ or $s(F)=2$.
Over such fields, every subset of $F^{\ast}/F^{\ast 2}$ is the value set of a unique anisotropic quadratic form over $F$ by [Lam05, Chapter XI. Theorem 7.19 (4)]. In particular, we have $\mathsf{D}_{F}(\pi_{U,0})=U$ and $\mathsf{D}_{F}(\pi_{U,1})=U\cup-U$. One now readily sees that each form in $\mathsf{P}_{n,m}(F)$ has exactly one preimage under $\varphi_{n,m}$.
Note that $\overline{C}$-fields are examples of so-called rigid fields whose Pfister number, an invariant related to the symbol length, was studied by the second author in [Lor23].
@@
Note that $p(F)$ is finite if some $d_{m}$ is finite. It is easy to calculate $s_{m}$ in terms of the level $s(F)$.
@@
  3. (c)
    Note that $p(F)\leq 2^{m}$ iff $q_{m+1}=0$ iff $q_{k}=0$ for all $k\geq m+1$. Thus, one readily sees that if $d_{m}<\infty$, then $p(F)\leq 2^{m+d_{m}+s_{m}}$.
@@
Let $0\leq m\leq n$ be integers and let $F$ be a field with $d_{m}<\infty$. Then
$|\mathsf{P}_{n,m}(F)|\leq 2^{(n-m)(d_{m}+s_{m}-n+m)}\cdot\alpha_{n-m},$
where $\alpha_{n-m}$ is defined as in Corollary 2.2.
@@
The rest of the section is now devoted to finding upper bounds for the $d_{m}$ in order to optimize the above bound for the number of Pfister forms. We will see in the next result that we can express $d_{m}$ in terms of $s_{m}$ and the $q_{k}$ for $k\leq m$.
@@
We now want to collect estimates for the invariants $q_{m}$. To do so we will use the following extension of a classical result from I. Kaplansky.
@@
Let $F$ be a field with $q(F)=2^{d}<\infty$ and let $s\geq 0$ be the integer with $2^{s}\leq p(F)<2^{s+1}$.
@@ 3. Bounding the Symbol Length
$\mathcal{B}=\mathcal{B}_{0}\supseteq\mathcal{B}_{1}\supseteq\mathcal{B}_{2}\supseteq\ldots$
@@
If we have $\pi_{j}\in\mathsf{P}_{n,m}(F)$ for some $m\in\operatorname{\mathbb{N}}$, the resulting Pfister forms are hyperbolic or will lie in $\mathsf{P}_{n,m^{\prime}}(F)$ for some $m^{\prime}\geq m$. We can clearly omit the hyperbolic Pfister forms in our representation and thus now consider any of the resulting Pfister forms lying in $\mathsf{P}_{n,m^{\prime}}(F)$ for some $m^{\prime}\geq m$
If we have $m^{\prime}=m$, we can repeat the above until the considered slot in any newly introduced Pfister form lies in $\mathcal{B}_{m}$. If we have $m^{\prime}>m$, we first substitute the representation of this form with an isometric Pfister form with $m^{\prime}$ slots equal to $-1$ and all other slots being products of $-1$ and of elements lying in $\mathcal{B}_{m^{\prime}}$. We can then apply the above procedure to this new form.
By repeating this process for all of the $\pi_{j},j\in\{1,\ldots,k\}$ and all slots that neither lie in the appropriate $\mathcal{B}_{m}$ (up to a factor $-1$) nor are equal to $-1$, and possibly rearranging the terms, we will eventually obtain a representation as desired. ∎
@@
If $F$ is nonreal of level $s(F)=2^{s}$, it is obviously enough to restrict the outer sum occuring in Proposition 3.4 to $m\geq\max\{0,n-s\}$.
@@
as in Proposition 3.4. Let $X\in\mathcal{P}_{m,r}$ and $Y\in\mathcal{P}_{m,r}^{\prime}$. Then all forms $\pi\in\mathsf{P}_{n,m}(F)\cup\mathsf{P}_{n,m+1}(F)$ that occur in this representation and that have the elements of $X\cup Y$ as slots (up to a factor $-1$) can be replaced by a single Pfister form using Lemma 3.3 just as in the proof of Corollary 3.7.
Conversely, let some $\pi=\pi_{U,m}$ from this representation be given and let $B=\{-b_{1},\ldots,-b_{n-m}\}$ be the set of slots $\neq-1$ of $\pi$. Let $r$ be maximal such that there is an $X\in\mathcal{P}_{m,r}$ with $X\subseteq-B$. We clearly have
@@
There is thus a subset $Y\subseteq\mathcal{A}_{m}^{\prime}$ with cardinality $|Y|=n-m-2r-1$ such that $X\cup Y\subseteq-B$.
Hence, for all Pfister forms of type $\pi_{U,m}$ for a suitable $U$, there is an $r\in\operatorname{\mathbb{N}}$, $X\in\mathcal{P}_{m,r}$ and $Y\in\mathcal{P}_{m,r}^{\prime}$ such that all slots except one are either $-1$ or lie (up to a factor $-1$) in $X\cup Y$. Therefore, our above strategy to replace certain Pfister forms by a single one can be applied to all forms on the right hand side in (1).
The symbol length $\operatorname{\mathrm{sl}}_{n}(\varphi)$ can thus be bounded by the cardinality of
@@
We now close the article with a comparison of these bounds for level $s(F)=2^{s}$ with $s\in\{0,1,2\}$. (Recall that conjecturally, if $s(F)<\infty$ then $q(F)<\infty$ implies $s(F)\leq 4$.) Of interest in this context are certain fields of iterated Laurent series that realize a given value $q(F)$ for the respective value $s(F)$, such as
  • $s=1$, $q=2^{d}$ ($d\geq 0$): $\operatorname{\mathbb{C}}(\!(t_{1})\!)\ldots(\!(t_{d})\!)$;
  • $s=2$, $q=2^{d}$ ($d\geq 1$): $\mathbb{F}_{3}(\!(t_{1})\!)\ldots(\!(t_{d-1})\!)$;
  • $s=4$, $q=2^{d}$ ($d\geq 3$): $\operatorname{\mathbb{Q}}_{2}(\!(t_{1})\!)\ldots(\!(t_{d-3})\!)$.
(Note that for any field $F$ with $s(F)=4$, one has $q(F)\geq 8$, see [EL73, Theorem 2.7].)
Before giving some explicit estimates, we want to recall the upper bounds that we found above.
First of all, we want to collect upper bounds for $d_{m}$ for some cases. In the following table, we have summarized the upper bounds we know for $d_{m}$ obtained in Corollary 2.13:
$m$ $s$ 0 1 2
0 $d$ $d-1$ $d-1$
1 00 $d-1$ $d-3$
2 00 00 $d-3$
For $s_{m}$ we have the following values according to Remark 2.8 (a):
$m$ $s$ 0 1 2
0 00 $1$ $1$
1 00 00 $1$
2 00 00 00